This record is intended to indicate maintenance on the station that may affect the readings from the instruments. They are in reverse chronological order (for no particular reason).
12 February 2019 - The calibration constants for the solar irradiance were modified to remove an offset in the preamplifier and to better indicate the correct value. Also switched to a better equation for the heat index. Other minor changes were made that don't affect the displayed values.
17 May 2018 - Made minor modifications to the front panels that are shown for the graphical current conditions from the web page. Also modified the "accumulate" time (when the values taken each second are accumulated into the archived values given once a minute). The time in the archived data will usually be shown at 30 seconds into the minutre rather than 22 seconds into the minutes as seems to have been normal in the past.
3 July 2013 - There is a connection problem near the computer with the solar radiation gauge. The wires have been fiddled to make a good connection - for now. Unfortunately, the fiddling caused the computer to register 0.09" rainfall that does NOT represent actual rainfall.
2 July 2013 - The anemometer and hygrometer were both replaced. The new devices are the same model as the previous devices but the age of those devices recommended that they be replaced to maintain performance.
25-27 June 2013 - On 25 June the hard disk in the computer crashed. Some data were lost for 25 June (may be recoverable with some effort). Replacement hardware had already been obtained but it was necessary to do some testing before it could be installed. The new hardware was brought online on 27 June 2013.
Second, the LabVIEW program was significantly rewritten. Initially this was to handle the newer USB device and switch to the DAQmx hardware drivers. Additional changes included changing the user interface to be event driven rather than polled. And there is now a configuration file that specifies most of the details of the program (like which hardware channel a particular signal is attached to, the sampling order of the channels, the calibration equation for each channel, etc.). The configuration file greatly improves the ability to modify hardware without requiring a program rebuild.
21 April 2009 - there has been a significant offset in the solar radiation gauge for a long time. The offset on the amplifier was adjusted with the sensor covered. It was not possible to get to 0 with no light - there is about 25 W/m^2 residual signal. It will be necessary to modify the amplifier to reduce the offset any further.
6 January 2005 - a heater was wrapped around the rain gauge to see if it would work to melt the snow and not cause other problems. The precipitation that was recorded for the January 6-7 actually fell starting about Tuesday night.
7 May 2004 - the cover of the rain gauge was taken off for inspection and the wind caused the bucket to tip, registering 0.01" rain when nothing actually fell.
3 May 2004 - for an unknown reason the rain gauge registered 0.01" rain when nothing actually fell. Suspect possibility of wind causing a bucket tip.
7 August 2002 - the programs for looking at past data on the web were changed to report dew point for all previous data. For dates before 1 August 2002 the dew point is calculated on the fly; for later dates the data from the original acquisition is used. It is no longer possible to view the wind chill factor for any dates.
25 July 2002 - the program was modified to save the dew point in the place of the wind chill factor (which is no longer being reported).
July 2002 - the computer keeps locking up at random times. All possible patches for Windows 2000 were installed and a new video driver. Doesn't seem to have made any difference.
7 March 2002 - the computer operating the instruments was replaced with a 1.0 GHz Pentium III. Response in providing the values to the web was improved by about a factor of 10. The extra processing power may also allow us to provide more enhancements to the display screen in the future.
8 February 2002 - the supply voltage on the wind direction was sufficiently low that it was often interpreted as being off and the direction defaulted to 290o. Some circuitry was modified and the voltage was increased to be sure that it wasn't misinterpreted. The indicated direction is fairly close to correct now.
20 November 2001 - The new humidity instrument was connected to the system. The relative humidity will no longer top out at about 80%.
20 November 2001 - The new wind direction indicator was connected to the system. The calibration is not exceptionally good (probably within 30o). Hopefully we can improve on that calibration soon.
1 November 2001 - The original temperature indicator was reattached.
30 October 2001 - As part of the upgrade the installers mistakenly disconnected the temperature indicator. This also causes the correction of the barometric pressure to mean sea level to be incorrect (it involves the temperature). The absolute pressure is still correct.
29 October 2001 - Several of the instruments are being upgraded. The wind direction indicator was disconnected in preparation for installing the new indicator during the afternoon.
12 June 2001 - The computer hung during the only rainstorm of value in the month of June (at least so far). The SWKT station recorded 0.12" of rain during the period that it was hung.
20 December 2000 - The zero level on the barometer was raised by 0.034 V
(0.20 in Hg) and the sea level correction algorithm was corrected - changed
from a strict additive constant to an exponential dependence:
Pmsl = Pstation * exp(0.03415*H/Tm)
where Pmsl is the mean sea level pressure, Pstation is the station pressure (absolute pressure at the station), H is the altitude of the station above mean sea level in meters, and Tm is the 12-hour mean temperature (average of the current temperature and the temperature 12 hours ago) in Kelvin.
17 August 2000 - The zero level on the barometer was lowered by another 0.021 V (0.13 in Hg). This exactly calibrates our barometer with both SWKT and aiport values at 30.16 in Hg and 84 F air temperature. The temperature sensitivity appears to be less pronounced but is still there.
17 August 2000 - The cable from the solar radiation sensor was changed to a shielded cable. Grounds were all reconnected to be sure we had good contact (the shield of the new cable, the tower, and the amplifier). This appears to have removed all the noise on that line.
20 July 2000 - The zero level on the barometer was lowered by 0.037 V (0.22 in Hg) to better match the SWKT and airport values. Tried various things to see where the temperature sensitivity of the barometer was with no luck.
28 June 2000, 8:30 am - during work on the solar insolation sensor a spurious 0.04" rain was recorded due to noise introduced on the lines
28 June 2000, 8:30 am - added filter capacitors to the solar insolation sensor to cure significant noise problems on that sensor (problems cropped up on 22 June as random noise spikes and became progressively worse)
19 June 2000, noon - solar insolation amplifier has been reading negative values for about 36 hours. Checked all aspects of the amplifier and can find nothing wrong. Re-seated the amplifier (741) in the socket to avoid possible contact corrosion problems.
18 May 2000 8:35 am - checked the calibration of the hygrometer - wet-bulb readings indicated 72% humidity while the sensor was reading 73%.
5 April 2000, 2:30 pm - adjusted the hygrometer to match the web-bulb reading taken on the roof near the sensor. The screw one the top of the sensor was turned clockwise to bring it from 52.5 to 25% reading. The support people at Texas Electronics indicated that this is probably good for a year after which the sensor will break.
19 February 2000 - station was put into operation.
Return to Weather Home Page.
Copyright © 2019. Brigham Young University. All rights reserved.